A small object has a mass of 3.0 × 10^-3 kg and a charge of -37C. It is placed at a certain spot where there is an electric field. When released, the object experiences an acceleration of 3.4 × 10^3 m/s2 in the direction of the +x axis. Determine the electric field, including sign, relative to the +x axis.
1
Expert's answer
2020-09-16T10:12:50-0400

Given

mass is "m= 3.0 \u00d7 10^{-3} kg"

Charge is "q=-37 C"

Acceleration is "a=3.4 \u00d7 10^3 m\/s^2"

According to Newton's second law

"F=ma"

And "F=qE"

From the above two equations,

"qE=ma"

The electric field strength is

"E=\frac{ma}{q}=\frac{( 3.0 × 10^{-3} kg)(3.4 × 10^3 m/s^2)}{-37 C}=-0.276 N/C"

Direction is along negative x axis.